OpenStudy (anonymous):

The solution of 6 divided by the quantity n squared minus 2 n minus 1 divided by n equals 1 divided by the quantity n minus 2 is n = 4. True False

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (anonymous):

6/n^2-2n - 1/n = 1/n-2 is n=4

OpenStudy (anonymous):

@Directrix

OpenStudy (blurbendy):

can you draw the equation

OpenStudy (anonymous):



OpenStudy (blurbendy):

if n = 4, then it is true

Directrix (directrix):

I agree with ^ @blurbendy n = 4

OpenStudy (anonymous):

thanks :)

Directrix (directrix):

I thought there might be some restrictions on n but I could not find any.